OpenStudy (anonymous):

How do I even start simplifying this expression? {[(x+4)^1/2]-3}/x-5 ? Its a limits continuation problem with x not equaling 5. What can I do to alter the denominator so that F is continuous at 5?

OpenStudy (rational):

are you trying to evaluate the limit as x->5 ?

OpenStudy (amistre64):

([(x+4)^1/2]-3)/x-5 0/0, try LHop?

OpenStudy (anonymous):

Yes im trying to evaluate the limit as x-->5. And Im not using l hop yet

OpenStudy (amistre64):

use a conjugate?

OpenStudy (amistre64):

a table method is to try to evaluate it for values of x closer and closer to x=5, if the lft and right are approaching the same value, then that the limit

OpenStudy (anonymous):

Conjugate for the numerator or the denominator? also i find it confusing how to use a conjugate if one of the numbers are not under the square root sign.

OpenStudy (anonymous):

using foil seems to jumble it up terribly

OpenStudy (anonymous):

I cant seem to find a way to make the denominator not equal 0

OpenStudy (amistre64):

by cheating with the wolf, this simplifies to 1/(sqrt(x+4)+3) which is the conjugate of the top

OpenStudy (amistre64):

(sqrt(x+4) -3)(sqrt(x+4)+3) = x+4-9 = x-5

OpenStudy (amistre64):

so conjugate was indeed the way to approach it

OpenStudy (amistre64):

recall (a-b)(a+b) = a^2 - b^2

OpenStudy (anonymous):

Thank you all very much!

OpenStudy (amistre64):

youre welcome :)